• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 16:11:30 on March12,2007

    Entry number 195712

    Followups:

    keyword=L-arm TC messed up, fix

    After the trigger download at 14:31 today the L-arm coinc time
    (D.ctimel) became again messed up -- out of range. See fig 1.
    But the R-arm is ok, and was always ok. So I adjusted the database
    and reloaded the trigger the official way. Run 1459 to check.
    Now it looks fine again -- see fig 2. Sorry for the confusion
    but at least I think the database is ok now.

    Also I would suggest we keep the EDTM pulser on at 5 Hz.
    It was apparently off for some runs but is on now.


    FIGURE 1

    FIGURE 2